OpenStudy (anonymous):

y varies jointly with p and q when p=1 and q=2 y=30 so I have to find y when p = 10 and q= 3 im pretty sure the answer is 300 but I want to make sure

OpenStudy (marissalovescats):

Joint variation is z=kxy.

OpenStudy (marissalovescats):

Or in your case, y=kpq where K is a constant number.

OpenStudy (marissalovescats):

Soo 30=k(1)(2)

OpenStudy (marissalovescats):

30=2k. What does K=? @johnrocks989

OpenStudy (marissalovescats):

Once we have found K, which we know is 15 because 30/2 is 15 We plug in p and q So y=(15)(10)(3) So what does y=?

OpenStudy (anonymous):

this math is hurting my head lol thank you for helping me I think I have it now :) marrisa

OpenStudy (marissalovescats):

You're welcome lol